OpenStudy (anonymous):

simplify cotx sinx-sin(pi/2-x)+cotx plz help i'm so lost

OpenStudy (anonymous):

sin(pi/2-x)=cos x

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

@surjithayer okay is that the formula i need to use?

OpenStudy (jdoe0001):

\(\bf cot(x)sin(x)-sin\left(\frac{\pi}{2}-x\right)+cot(x)\\ \quad \\\implies cot(x)sin(x)-cos(x)+cot(x)\\ \quad \\ \cfrac{cos(x)}{sin(x)}sin(x)-cos(x)+\cfrac{cos(x)}{sin(x)}\)

OpenStudy (jdoe0001):

recall that \(\bf cot(\theta) = \cfrac{cos(\theta)}{sin(\theta)}\)

OpenStudy (anonymous):

@jdoe0001 yes i do now how exactly do i plug the question in?

OpenStudy (jdoe0001):

well, you just want to simplify I gather, so cancel any like-terms and add, see what you're left with

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(x)}{\cancel{sin(x)}}\cancel{sin(x)}-cos(x)+\cfrac{cos(x)}{sin(x)}\)

OpenStudy (anonymous):

@jdoe0001 yup so would my answer be cosx or 2 cosx

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(x)}{\cancel{sin(x)}}\cancel{sin(x)}-cos(x)+\cfrac{cos(x)}{sin(x)}\implies cos(x)-cos(x)+\cfrac{cos(x)}{sin(x)}\) see anyting you can cancel out?

OpenStudy (anonymous):

@jdoe0001 yes -cos(x) and cos(x)

OpenStudy (jdoe0001):

yeap, and you're left with that, with will be just cot(x) :)

OpenStudy (anonymous):

omg thank you so much! @jdoe0001

OpenStudy (jdoe0001):

yw